Tyrion:

Show all work to multiply quantity 2 plus the square root of negative 25 end quantity times quantity 4 minus the square root of negative 100 end quantity

Tyrion:

@snowflake0531

Tyrion:

@itsmehjay

snowflake0531:

\[am~I~righ t ~ that~this~is~ (2+ \sqrt{-25})(4-\sqrt{-100})\]

Tyrion:

yes

snowflake0531:

Do you know the FOIL method? \[(a+b)(c+d) = ac+ad+bc+bd\]

Tyrion:

no maam

snowflake0531:

Although the second one is minus sqrt-100, you can take it as plus, negative sqrt-100

snowflake0531:

@tyrion wrote:
no maam
welp, now you know it now

snowflake0531:

So can you use that method and simplify that ^^^^

snowflake0531:

Or do you want me to write this out for you?

Tyrion:

i never used the method so idek how to take the numbers and put them in their spot

snowflake0531:

\[(2+\sqrt(-25))(4+(-\sqrt(-100))\] \[2(4)+2(-\sqrt{-100}) + (\sqrt{-25})4 + (\sqrt{-25})(-\sqrt{-100})\] Do you know how to simplify that?

Tyrion:

no lie,,ion kno wat goin on,,this is a pre test question,,we aint learn dis yet

snowflake0531:

actually wait, I did it hte wrong way shoot shoot shoot

snowflake0531:

We have to pull out imaginary numbers do you know what that is?

Tyrion:

not at all

snowflake0531:

We have \[(2+\sqrt{-25})(4-\sqrt{-100})\] \[imaginary~number~is~i~and~i~is~ \sqrt-1\]

snowflake0531:

And because we have that i is sqrt-1 \[(2+i \sqrt25)(4-i \sqrt100)\]

snowflake0531:

From there, we have \[(2+5i)(4-10i)\]

snowflake0531:

Do you get it so far?

Tyrion:

yuh

snowflake0531:

Okay so now we use the FOIL method \[2(4) + 2(-10i) + 5i(4) + (5i)(-10i)\]

snowflake0531:

This simplifies to\[8-20i+20i+50\]

snowflake0531:

which then equals \[58\]

Tyrion:

ok,,ima tryna write dis stuff down

snowflake0531:

Okay well, dm me if you have questions, since right now, I have to go

Tyrion:

ok thnx

snowflake0531:

yw~

surjithayer:

\[(2+5\iota)(4-10\iota)\]\[=2(2+5\iota)(2-5\iota)\]\[=2[(2)^2-(5\iota)^2]\]\[=2[4-25\iota^2]\]\[=2(4-25(-1))\] \[=2(4+25)\] \[=2 \times 29\] =58

kittybasil:

tyrion wrote:
Show all work to multiply quantity 2 plus the square root of negative 25 end quantity times quantity 4 minus the square root of negative 100 end quantity
Okay, let's first turn the words into numbers. "Quantity 2" = 2 "square root of negative 25" = \(\sqrt{-25}\) Note: Look at the negative in there. This means you will have an imaginary number "\(i\)" because \(i=\sqrt{-1}\)) "times quantity 4" means multiply by 4, or \(\times4\) "minus the square root of negative 100" = \(\sqrt{-100}\) Note: again, this will have an imaginary number because it has a negative inside the square root. You will notice I mentioned imaginary numbers twice. This only happens when you have negatives inside a square root. What you need to do to resolve that negative is isolate/remove it from the rest of the numbers inside the square root. Here is a step by step example:\[\sqrt{-x^2}=\sqrt{(-1)(x^2)}\]\[=\sqrt{-1}\cdot\sqrt{x^2}\]\[=i\cdot x\]As previously mentioned, the indicator "\(i\)" means the square root of NEGATIVE ONE. Not a random negative number. This is why you have to separate out the negative from the rest of the numbers.

kittybasil:

Sorry for the formatting error. I would redo it but that's a lot of work... let me know if you get everything! Continuing on though, we must put everything together. Your end result should be this:\[(2+\sqrt{-25})(4-\sqrt{-110})\]Normally I'd just leave the next steps up to you. However, since it's your first time working with imaginary numbers, I'll do that work to save you some time:\[(2+i\sqrt{25})(4-i\sqrt{110})\]Now the negative square root part has been resolved. Where should we go from here? Share your thoughts @tyrion